Assume that a consumer purchases a combination of products A and B. The MUa is 5 and the Pa is $5. The MUb is 4 and the Pb is $6. What should this consumer do to maximize utility?
Answer #1

Answer

In order to maximize utility a consumer should consume that quantity of A and B such that (MUa/Pa) = (MUb/Pb). Here we have (MUa/Pa) = 5/5 = 1 and (MUb/Pb) = 4/6 = 0.67. Hence, (MUa/Pa) is not equal to (MUb/Pb)

Thus he is not maximizing utility. According to law of diminishing marginal utility As consumer consumes more unit, Marginal utility derived from additional utility must decline.

So in order to make MUa/Pa equal to MUb/Pb. We have to increase consumption of A and decrease consumption of B, which will result in decrease in MUa and increase in MUb and this will continue till the point comes where MUa/Pa = MUb/Pb.

Hence Consumer should increase consumption of Product A and decrease consumption of B.